The subraction of a matrix B may be considered as the addition of the matrix(-1)B.Does the commutative law of addition permit us to state that A-B=B-A?If not,how would you correct the statement?


1
Expert's answer
2021-10-25T17:43:46-0400

If the subtraction A - B is considered a the addition A + (-B), then the commutative law should be stated as follows :

A + (-B) = (-B) + A.

That is, A - B = -B + A.

Thus, the corrected statement is A - B = -B + A, not B - A.